[Date Prev][Date Next][Thread Prev][Thread Next][Date Index][Thread Index]

Mars Pathfinder Update - September 14, 1997

[Checkout the latest Mars Pathfinder rover images of the rocks Half Dome 
 and Moe:



 Ron Baalke

14 September 1997, 8:45 p.m. PDT

The lander and rover woke up today at 1:00 p.m. PDT.

The Sojourner batteries have been expended and the rover
is now dependent on solar power. All data indicates that
both the lander and rover are in good health.

Due to an uplink error last week we did not run the
standard science sequence today. But we were able to run
a subset of the originally planed imaging by sending
real time commands from the Goldstone Deep Space Network
station. We acquired super resolution images and magnet
target data as well as a reduced set of ASI/MET data
which indicated two small dust devils passing over the
lander. The rover had no problems and executed all of
its command successfully.

Tomorrow the rover will traverse towards the rock Chimp
and we will continue to acquire ASI/MET and imaging data
from the lander.

For further information on the Mars Pathfinder Mission,
please call our Mission Status Report line at